# Are Large Language Models Memorizing Bug Benchmarks?

### v1

@@ Abstract
In this paper, we systematically evaluate popular LLMs to assess their susceptibility to data leakage from widely used bug benchmarks. To identify potential leakage, we use multiple metrics, including a study of benchmark membership within commonly used training datasets, as well as analyses of negative log-likelihood and n-gram accuracy. Our findings show that certain models, in particular codegen-multi, exhibit significant evidence of memorization in widely used benchmarks like Defects4J, while newer models trained on larger datasets like LLaMa 3.1 exhibit limited signs of leakage. These results highlight the need for careful benchmark selection and the adoption of robust metrics to adequately assess models capabilities.
@@ I Introduction
We use multiple metrics to detect potential leakage. First, we investigate whether benchmark data has membership within TheStack, a widely-used used pretraining code dataset. Following this, we apply two core metrics for leakage from prior work [9, 10]: Negative Log-Likelihood (NLL) and n-gram accuracy. NLL provides insight into model familiarity with code snippets; $N$-gram accuracy assesses the model’s ability to reproduce exact sequences. We apply these metrics to both well-known bug benchmarks, and a new dataset of high-quality code repositories from 2024, which we mined from GitHub. This new dataset is less likely to have appeared in models’ training, which allows us to compare model performance between potentially familiar data versus likely novel code.
Our findings suggest that older models, in particular codegen-multi, exhibit very high $n$-gram accuracy and low NLL on benchmark data, indicating a higher likelihood of memorization. Our evidence suggests that newer models trained on more extensive datasets, like LLaMa 3.1, show less memorization. Nonetheless, across all metrics, models, and families, Defects4J — arguably the most widely-used bug benchmark — consistently exhibits the highest rate of potential memorization. These results underscore the importance of carefully selecting benchmarks to ensure reliable evaluations.
@@ II Methodology
We select widely used bug benchmarks across common programming languages, gathering ground-truth files containing reference solutions for each bug fix. To provide a likely not-leaked dataset for comparison, we also curated a set of recent open-source repositories from GitHub.
@@
We also collected a new dataset of 3214 GitHub repositories written in Java and Python. We targeted repositories from 2024 to reduce the likelihood that current state-of-the-art LLMs have seen them. To narrow our selection to likely high-quality, indicative repositories, we applied a minimum threshold of 100 stars, a proxy for community interest and engagement. To ensure that we were mining novel datasets for LLMs, we applied the MinHash technique [16] with Locality Sensitive Hashing (LSH) [17] to filter repositories potentially overlapping with existing models’ training data. Because of code duplication, some files in new 2024 repositories might be identical to those in older repositories. As a result, we collected all repositories created between 2022 and the end of 2023 with greater than 100 stars to exclude the repositories from 2024 that contained data from these older, highly popular repositories.
Finally, to manage computational resources, we randomly sampled files from the collected data (k=250) for our evaluation.
@@
We select a combination of models used for fault localization [25], program repair [26], and vulnerability detection [27]. Table II shows model information. These models are from families of well-known code-related models, including codegen-multi, the LLaMa 3.1 family, the Gemma 2 family, StarCoder, and Mistral. Note, we could not find detailed information about the training data for Mistral’s models, apart from anecdotal evidence on public forums.
@@ III Results
This section presents results assessing possible leakage of bug benchmarks in base models, using the metrics described in Section II-C: membership in TheStack (Section III-A), Negative Log Likelihood (Section III-B), and $N$-gram accuracy (Section III-C). We also perform a regression analysis of model characteristics, NLL, and $N$-gram accuracy (Section III-D) to better understand characteristics of models that influence data leakage. (Note that we subsequently discuss implications in Section IV, and limitations and threats to the validity of our experiments in Section V.)
@@ III-B Negative Log Likelihood
Interestingly, codegen-multi 6B exhibits low NLL (0.38) on SWEBench-Lite compared to other models, despite being the oldest model in our evaluation, trained on older data, and the fact that SWEBench-Lite was published recently. This is because the projects in the benchmark existed prior to benchmark publication, as we also saw in the membership analysis (Table III). Moreover, although SWEBench-Lite is a new benchmark, the bug fixes date as early as 2017.
@@
III-C 5-Gram Match
Figure 5: 5-gram accuracy by model and dataset. For each model family (see Section II-B), we report 5-gram accuracy of either the only model available (e.g., StarCoder) or the biggest model in the family (e.g., LLaMa 3.1 70B).
Figure 5 shows $5$-gram match results, a complementary assessment of potential model memorization (full tables, including other models are in the appendix). Note that, due to differences in model vocabularies and tokenization, interpretation of $5$-gram match (similar to NLL) can differ across models.
Defects4J consistently exhibits the highest $5$-gram match across all model families. Conversely, $5$-gram accuracy on Gitbug-Java is relatively similar to that of new repositories for all models, which aligns with the expectation that most repositories in the Gitbug-Java benchmark were not included in pretraining data (as detailed in Table I). For example, both codegen-multi and Gemma 2 show significantly higher $5$-gram match differences between Defects4J and new repositories (i.e., $34$ percentage points and $14$ percentage points, respectively). Moreover, codegen-multi achieves over $82\%$ $5$-gram match on Defects4J, strongly suggesting that it has likely memorized much of the benchmark’s solutions.
As expected, new repositories generally exhibit lower $5$-gram match across all models, with averages of 47% and 48% for Java and Python, respectively. These values happen likely due to the presence of common coding patterns [29]. Conversely, and in line with our NLL findings, codegen-multi shows a notably high $5$-gram match on SWEBench-Lite, even though it is a recently published dataset, as it incorporates older data.
@@ III-D How Do Model Characteristics Influence Risk of Leakage?
To gain deeper insights, and using the data collected during the study, we estimate regression models for negative log-likelihood (NLL) and 5-gram accuracy. The regressions allow us to explore relationships and (potentially) causality that metrics alone cannot reveal. For example, while codegen-multi presents low NLL values and high 5-gram accuracy, it is unclear what drives this outcome. By employing regression analysis, we can identify factors such as training parameters and budget as significant influences on these metrics. Therefore, for NLL, we use a mixed-effects linear model, while for 5-gram accuracy (i.e., a probability), we use a mixed-effects logistic regression. As predictors, we include the model’s number of parameters and number of tokens used during pretraining (i.e., training budget), which serves as a proxy for the unique token count. We acknowledge that reporting the exact number of unique tokens would be a more precise metric, but such data is often unavailable. We also account for potential variability caused by differences in the datasets by including dataset type as random effect.
Table IV shows results. For NLL, the intercept represents the average NLL value expected when all predictions are zero (1.497). The results suggest that increasing the training budget by one trillion tokens raises NLL by 0.091 units on solution files, while the effects of model parameters are negligible. For 5-gram accuracy, the intercept represents the log-odds when all predictors are zero, translating to a probability of approximately 63% (i.e., $e^{-0.457}$). This suggests that, by default, the models achieve 63% accuracy in predicting 5-grams. While this may initially seem like a high value, it’s actually not surprising given that code is highly repetitive and follows natural patterns. The positive coefficient for parameters (0.003) indicates that increasing the number of parameters leads to higher 5-gram accuracy. Conversely, the training budget shows a highly significant negative effect (-0.035), suggesting that the likelihood of achieving higher 5-gram accuracy decreases (and, consequently, memorization) as the training budget increases.
@@
Concretely, codegen-multi has the smallest known training budget of all models evaluated, with approximately 0.5T tokens. Due to its smaller budget, the resulting model generalizes less across all seen code and is more susceptible to memorization, as seen in Figure 1. The high $5$-gram match and low NLL support this intuition, especially for Defects4J. Conversely, newer models, like StarCoder, exhibit much lower memorization of Defects4J, despite its existence in all versions of TheStack. As StarCoder’s training budget is significantly larger than codegen-multi’s (3.5T tokens), Defects4J constitutes a smaller percentage of the overall training data. Regression results also imply that models with more parameters tend to exhibit higher n-gram accuracy and, consequently, memorize more. For example, Figure 6(b) shows that LLaMa 3.1 70B accurately predicts strings, class names, and if-statements in contexts where these predictions might not be immediately apparent, which may indicate memorization. Conversely, while the 8B (Figure 6(a)) version predicts the patch more accurately, it fails to identify the subsequent classes.
@@
Intercept 1.497 (0.251) *** -0.457 (0.020) ***
Parameters $-6.2e^{-4}$ (0.003) *** 0.003 ($7.0e^{-4}$) ***
Training budget 0.091 (0.013) *** -0.035 (0.003) ***
Note: *** p-value < 0.001
@@ IV Discussion
Our evaluation provides compelling evidence that data leakage is an especially significant issue for Defects4J (V1.5).This is evident from the lower NLL values, higher 5-gram accuracy. In particular, codegen-multi achieves $82\%$ 5-gram accuracy on Defects4J, while both CodeLlama 7B and Gemma-2 27B attain $64\%$. Moreover, given that Defects4J is incorporated into the widely-used pretraining dataset (TheStack), with $80\%$ membership, eliminating this leakage in future models is likely nearly impossible. We also observe similarly low NLL values in SWEBench-Lite for codegen-multi.
Newer benchmarks BugsInPy and BugsCpp exhibit lower leakage risk. A smaller percentage of their repositories are indexed in the latest version of TheStack. While BugsInPy and BugsCpp exhibit slightly lower NLL compared to new repositories, their 5-gram accuracy and NLL values are not significantly different from newer, more-likely-unseen benchmarks like Gitbug-Java. This suggests that, for now, researchers can use these benchmarks with a relatively low risk of data leakage. However, as pretraining datasets continue to evolve, we recommend that researchers regularly assess TheStack membership, $5$-gram match and NLL values of these benchmarks, especially compared to more recent data, to monitor and mitigate potential data contamination.
We also observe that newer models, e.g., the LLaMa 3.1 family, seem to have lower leakage (likely due to being trained on more data). Nonetheless, we still observed cases where LLaMa 3.1 can output solution files despite little context (e.g., Figure 6(a)).
@@ V Limitations and Threats
Data Collection: Despite efforts to filter out older GitHub repositories when collecting new data, we anecdotally observed instances where files appeared to be adaptations of existing files (e.g., from 2018555https://bit.ly/3Od3nC6.) Our filtering process may not have perfectly excluded legacy code. We moreover cannot guarantee that the new repositories are identically distributed compared to those in the benchmarks. To mitigate this issue, given that the repositories in the benchmarks tend to be highly recognizable projects, we applied a >100-star filter to argue that the selected projects are of comparable quality.
Train + Test Splits: For benchmarks like Defects4J (v1.5) and BugsInPy, the patch files and buggy files are very similar (typically a patch involves only changing a small number of lines). LLMs may have only seen the train split of these benchmarks at pretraining time. This would result in high 5-gram match and low NLL on patch files, even if only buggy files were leaked. Nonetheless, we observed multiple cases where models output patched files verbatim. We mitigate this by looking at trends in NLL and 5-gram match rather than absolute numbers in our analysis.
Forgetting: The findings presented in this paper primarily address leakage in the context of base models. Empirical results show that models can “forget" portions of their pretraining data during fine-tuning [30]. That said, while full-scale model fine-tuning may reduce leakage risks, more recent fine-tuning strategies—such as the addition of adapter layers—often "freeze" pretrained weights. This practice can inadvertently increase the likelihood of data leakage. Furthermore, larger models exhibit a stronger tendency to memorize training data [31], suggesting that while fine-tuning may mitigate leakage for smaller models, it is more likely to exacerbate leakage concerns in larger models. We leave empirical assessment of these risks to future work.
@@ VI Related Work
Large Language Models (LLM): LLMs have shown promise across a wide range of natural language [32, 33] and code generation tasks [34, 35, 36]. Current state of the art LLMs include open source models such as CodeGen, CodeLlama, LlaMa 3.1, Gemma and PhiCoder and closed-source models such as GPT-4o, Claude 3.5, and Gemini 1.5. LLMs have shown state of the art performance in APR, suggesting patches to buggy code segments, detecting vulnerabilities and helping pinpoint buggy lines in a piece of code [25, 26, 37].
@@
Measuring Data Leakage: This task is still an open challenge. Three common ways to measure data leakage are perplexity or NLL, n-gram match, and model performance. Perplexity approaches  [10, 9] quantify leakage by comparing loss across benchmarks. N-gram approaches [9, 38], prompt a model with context and measure the overlap of generated code with the ground truth. Finally, performance approaches [39, 40] measure the difference in performance between a benchmark and a transformed version of the benchmark. While we use techniques from the literature, we are the first to apply these metrics to bug benchmarks widely used in the software engineering community. We avoid measuring performance with transformations, as base models perform very poorly on untransformed bug benchmarks without fine-tuning (e.g.: we experimented with LLaMa 3.1 70B and achieved nearly 0% performance at fault localization on Defects4J).

### v2

@@ Abstract
In this paper, we systematically evaluate popular LLMs to assess their susceptibility to data leakage from widely used bug benchmarks. To identify potential leakage, we use multiple metrics, including a study of benchmark membership within commonly used training datasets, as well as analyses of negative log-likelihood and 5-gram accuracy. Our findings show that certain models, in particular codegen-multi, exhibit significant evidence of memorization in widely used benchmarks like Defects4J, while newer models trained on larger datasets like LLaMa 3.1 exhibit limited signs of leakage. These results highlight the need for careful benchmark selection and the adoption of robust metrics to adequately assess models capabilities.
@@ I Introduction
We use multiple metrics to detect potential leakage. First, we investigate whether benchmark data has membership within TheStack, a widely-used used pretraining code dataset. Following this, we apply two core metrics for leakage from prior work [9, 10]: Negative Log-Likelihood (NLL) and 5-gram accuracy. NLL provides insight into model familiarity with code snippets; 5-gram accuracy assesses the model’s ability to reproduce exact sequences. We apply these metrics to both well-known bug benchmarks, and a new dataset of high-quality code repositories from 2024, which we mined from GitHub. This new dataset is less likely to have appeared in models’ training, which allows us to compare model performance between potentially familiar data versus likely novel code.
Our findings suggest that older models, in particular codegen-multi, exhibit very high 5-gram accuracy and low NLL on benchmark data, indicating a higher likelihood of memorization. Our evidence suggests that newer models trained on more extensive datasets, like LLaMa 3.1, show less memorization. Nonetheless, across all metrics, and models, Defects4J — arguably the most widely-used bug benchmark — consistently exhibits the highest rate of potential memorization. These results underscore the importance of carefully selecting benchmarks to ensure reliable evaluations.
@@ II Methodology
We select widely used bug benchmarks across common programming languages, gathering ground-truth files containing reference solutions for each bug fix. To provide a likely not-leaked datasets for comparison, we also curated a set of recent open-source repositories from GitHub.
@@
We also collected a new dataset of 3214 GitHub repositories written in Java and Python. We targeted repositories from 2024 to reduce the likelihood that current state-of-the-art LLMs have seen them. To narrow our selection to likely high-quality, indicative repositories, we applied a minimum threshold of 100 stars, a proxy for community interest and engagement.
To ensure that we were mining novel datasets for LLMs, we applied the MinHash technique [16] with Locality Sensitive Hashing (LSH) [17] to filter repositories potentially overlapping with existing models’ training data. Because of code duplication, some files in new 2024 repositories might be identical to those in older repositories. As a result, we collected all repositories created between 2022 and the end of 2023 with greater than 100 stars to exclude the repositories from 2024 that contained data from these older, highly popular repositories. Finally, to manage computational resources, we randomly sampled files from the collected repositories ($250$ files per language) for our evaluation.
@@
We select a combination of models used for fault localization [25], program repair [26], and vulnerability detection [27]. Table II shows model information. [^1] We could not find detailed information about the training data for Mistral’s models, apart from anecdotal evidence on public forums. These models are from families of well-known code-related models, including codegen-multi, the LLaMa 3.1 family, the Gemma 2 family, StarCoder, and Mistral.
Note that we focus on open-source base models because our method requires computing the negative log-likelihood (NLL) on sequences, which is generally not possible with closed-source models. We exclude instruction-tuned models and concentrate solely on pretrained models before fine-tuning. Since instruction-tuned models are optimized for conversational formats, n-gram accuracy may be a less suitable metric for measuring memorization in these models.
@@ III Results
This section presents results assessing possible leakage of bug benchmarks in base models, using the metrics described in Section II-C: membership in TheStack (Section III-A), Negative Log Likelihood (Section III-B), and 5-gram accuracy (Section III-C). We also perform a regression analysis of model characteristics, NLL, and 5-gram accuracy (Section III-D) to better understand characteristics of models that influence data leakage. (Note that we subsequently discuss implications in Section IV, and limitations and threats to the validity of our experiments in Section V.)
@@ III-B Negative Log Likelihood
Interestingly, codegen-multi 6B exhibits low NLL (0.38) on SWEBench-Lite compared to other benchmarks and new data, despite being the oldest model in our evaluation, trained on older data, and the fact that SWEBench-Lite was published recently. This is because the projects in the benchmark existed prior to benchmark publication, as we also saw in the membership analysis (Table III). Moreover, although SWEBench-Lite is a new benchmark, the bug fixes date as early as 2017.
@@
III-C 5-Gram Accuracy
Figure 5: 5-gram accuracy by model and dataset. Due to space constraints, we selected a sample of the most relevant models. 5-gram accuracy for other models are consistent with the results displayed.
Figure 5 shows 5-gram accuracy results, a complementary assessment of potential model memorization (full tables, including other models are in the appendix). Note that, due to differences in model vocabularies and tokenization, interpretation of 5-gram accuracy can differ across models.
Defects4J consistently exhibits the highest 5-gram accuracy across all model families. Conversely, 5-gram accuracy on GitBug-Java is relatively similar to that of new repositories for all models, which aligns with the expectation that most repositories in the GitBug-Java benchmark were not included in pretraining data (as detailed in Table I). For example, both codegen-multi and Gemma 2 show significantly higher $5$-gram match differences between Defects4J and new repositories (i.e., $34$ percentage points and $14$ percentage points, respectively). Moreover, codegen-multi achieves $82\%$ 5-gram accuracy on Defects4J, strongly suggesting that it has likely memorized much of the benchmark’s solutions.
As expected, new repositories generally exhibit lower 5-gram accuracy across all models, with averages of 47% and 48% for Java and Python, respectively. These values happen likely due to the presence of common coding patterns [29]. Conversely, and in line with our NLL findings, codegen-multi shows a notably high 5-gram accuracy on SWEBench-Lite, even though it is a recently published dataset, as it incorporates older data.
@@ III-D How Do Model Characteristics Influence Risk of Leakage?
To gain deeper insights, and using the data collected during the study, we estimate regression models for the average values of NLL and 5-gram accuracy. The regressions allow us to explore relationships that metrics alone cannot reveal. Using regression analysis, we can identify factors such as training parameters and budget as significant influences on these metrics. Therefore, we use a simple mixed-effects linear model to predict the average NLL and 5-gram accuracy. As predictors, we include the models’ number of parameters and number of tokens used during pretraining (i.e., training budget), which serves as a proxy for the unique token count. We acknowledge that reporting the exact number of unique tokens would be a more precise metric, but such data is often unavailable. We also account for potential variability caused by differences in the datasets and tokenizers by including dataset and tokenizer type as random effect. Table IV shows results.
In a linear regression, the intercept represents the predicted value of the dependent variable (averages of NLL and 5-gram accuracy) when all predictors are at their reference values. To make the interpretation easier, we centered the predictors around codegen-multi values. Therefore, the reference value of parameters is 6B and the reference value of training budget is 1T tokens. The regression coefficients represent the change in the dependent variable for a one-unit increase in the predictor, holding other predictors constant.
For NLL, the predicted average is 0.744 when at reference level, i.e., parameters = 6B and training budget = 1T. For every 1B increase in number of parameters (above 6B), the predicted NLL average decreases by $0.002$ units (while holding training budget constant, at 1T). Similarly, for every 1T tokens increase in training budget, predicted NLL decreases by $0.014$ units. For 5-gram accuracy, the predicted average is 0.465 when predictors are at reference level. For every 1B increase in number of parameters, the predicted 5-gram accuracy increases by $0.001$ units. Similarly, a 1-unit increase in the training budget leads to an increase of $0.006$ in 5-gram accuracy. All reported coefficients have statistical significance.
Regression results reveal consistent trends across model families (e.g., LLaMa 3.1 or Gemma). For example, both LLaMa 3.1 8B and 70B were trained on the same data with a training budget of 15T tokens. However, the 70B model is approximately ten times larger than the 8B model, leading to an overall increase in 5-gram accuracy across all benchmarks (as shown in Figure 5) and decrease in NLL. Similarly, within the Gemma 2 family, the 2B model was trained on 3T tokens, while the 27B was trained on 12T. Here, we observe the same trend: average 5-gram accuracy increases in the 27B model and NLL decreases.
@@
Regression results also imply that models with more parameters tend to exhibit higher n-gram accuracy and, consequently, memorize more. For example, Figure 6(b) shows that LLaMa 3.1 70B, which has the same training budget as LLaMa 3.1 8B, accurately predicts strings, class names, and if-statements in contexts where these predictions might not be immediately apparent, which may indicate memorization.
@@
Intercept 0.744 (0.094) *** 0.465 (0.049) ***
Parameters -0.002 (0.001) *** $0.001$ ($4.0e{-4}$) ***
Training budget -0.014 (0.004) ** $0.006$ ($2.4e{-3}$) ***
Note: *** p-value < 0.001, ** p-value < 0.01, * p-value < 0.05
@@ IV Discussion
Our evaluation provides compelling evidence that data leakage is an especially significant issue for Defects4J (V1.5). This is evident from the lower NLL values and higher 5-gram accuracy. In particular, codegen-multi achieves $82\%$ 5-gram accuracy on Defects4J, while both CodeLlama 7B and Gemma-2 27B attain $64\%$. Moreover, given that Defects4J is incorporated into the widely-used pretraining dataset (TheStack), with $80\%$ membership, eliminating this leakage in future models is likely nearly impossible. We also observe similarly low NLL values in SWEBench-Lite for codegen-multi.
Newer benchmarks BugsInPy and BugsCpp exhibit lower leakage risk in almost all models. A smaller percentage of their repositories are indexed in the latest version of TheStack. While BugsInPy and BugsCpp exhibit slightly lower NLL compared to new repositories, their 5-gram accuracy and NLL values are not significantly different from newer, more-likely-unseen benchmarks like GitBug-Java. This suggests that, for now, researchers can use these benchmarks with a relatively low risk of data leakage. However, as pretraining datasets continue to evolve, we recommend that researchers regularly assess TheStack membership, 5-gram accuracy and NLL values of these benchmarks, especially compared to more recent data, to monitor and mitigate potential data contamination.
We also observe that the LLaMa 3.1 family seems to exhibit lower memorization of benchmark solutions. Nonetheless, we still observed cases where LLaMa 3.1 70B outputs solution files despite little context (e.g., Figure 6).
@@ V Limitations and Threats
Data Collection: Despite efforts to filter out older GitHub repositories when collecting new data, we anecdotally observed instances where files appeared to be adaptations of existing files (e.g., from 2018666https://bit.ly/3Od3nC6.) Our filtering process may not have perfectly excluded legacy code. We moreover cannot guarantee that the new repositories are identically distributed compared to those in the benchmarks. To mitigate this issue, given that the repositories in the benchmarks tend to be highly recognizable projects, we applied a >100-star filter to argue that the selected projects are of comparable quality.
Train + Test Splits: For benchmarks like Defects4J (V1.5) and BugsInPy, the patch files and buggy files are very similar (typically a patch involves only changing a small number of lines). LLMs may have only seen the train split of these benchmarks at pretraining time. This would result in high 5-gram accuracy and low NLL on patch files, even if only buggy files were leaked. Nonetheless, we observed multiple cases where models output patched files verbatim. We mitigate this by looking at trends in NLL and 5-gram accuracy rather than absolute numbers in our analysis.
Forgetting: The findings presented in this paper primarily address leakage in the context of base models. Empirical results show that models can “forget” portions of their pretraining data during fine-tuning [30]. That said, while full-scale model fine-tuning may reduce leakage risks, more recent fine-tuning strategies—such as the addition of adapter layers—often “freeze” pretrained weights. This practice can inadvertently increase the likelihood of data leakage. Furthermore, larger models exhibit a stronger tendency to memorize training data [31], suggesting that while fine-tuning may mitigate leakage for smaller models, it is more likely to exacerbate leakage concerns in larger models. We leave empirical assessment of these risks to future work.
@@ VI Related Work
Large Language Models (LLM): LLMs have shown promise across a wide range of natural language [32, 33] and code generation tasks [34, 35, 36]. Current state of the art LLMs include open source models such as codegen-multi, CodeLLaMa, LLaMa 3.1, Gemma 2 and and closed-source models such as GPT-4o, Claude 3.5, and Gemini 1.5. LLMs have shown state of the art performance in APR, suggesting patches to buggy code segments, detecting vulnerabilities and helping pinpoint buggy lines in a piece of code [25, 26, 37].
@@
Measuring Data Leakage: This task is still an open challenge. Three common ways to measure data leakage are perplexity or NLL, n-gram accuracy, and model performance. NLL based approaches  [10, 9] quantify leakage by comparing loss across benchmarks. n-gram accuracy approaches [9, 38], prompt a model with context and measure the overlap of generated code with the ground truth. Finally, performance approaches [39, 40] measure the difference in performance between a benchmark and a transformed version of the benchmark. While we build on established techniques from the literature, our work is the first to apply these metrics to bug benchmarks widely recognized in the software engineering community. Additionally, we conduct a comprehensive study of both recent models, such as the LLaMa 3.1 family, and earlier models like codegen-multi.